a1=5,

a2=15,

a3=125,

a4=1235,

a5=?


1
Expert's answer
2020-04-11T16:31:18-0400

"a_{n+1}=12\ldots n \,5", "n \geq 1," "a_1=5" for "n=0".

So "a_5=a_{4+1}=12345"

Answer: "12345".